# Continuous Domain Generalization

### v1

@@ Abstract
Real-world data distributions often shift continuously across multiple latent factors such as time, geography, and socioeconomic context. However, existing domain generalization approaches typically treat domains as discrete or evolving along a single axis (e.g., time), which fails to capture the complex, multi-dimensional nature of real-world variation. This paper introduces the task of Continuous Domain Generalization (CDG), which aims to generalize predictive models to unseen domains defined by arbitrary combinations of continuous variation descriptors. We present a principled framework grounded in geometric and algebraic theory, showing that optimal model parameters across domains lie on a low-dimensional manifold. To model this structure, we propose a Neural Lie Transport Operator (NeuralLTO), which enables structured parameter transitions by enforcing geometric continuity and algebraic consistency. To handle noisy or incomplete domain descriptors, we introduce a gating mechanism to suppress irrelevant dimensions and a local chart-based strategy for robust generalization. Extensive experiments on synthetic and real-world datasets—including remote sensing, scientific documents, and traffic forecasting—demonstrate that our method significantly outperforms existing baselines in generalization accuracy and robustness under descriptor imperfections.
@@ 1 Introduction
Distribution shift refers to the change in data distributions between training and deployment environments, where the input-output relationship learned during training no longer holds at test time. This discrepancy fundamentally compromises the reliability of learned models and motivates the study of Domain Generalization (DG) [52; 49; 23; 14; 43], which aims to train models on multiple source domains that generalize to unseen domains without access to their training data. While early DG methods treat domains as independent entities, recent work recognizes that distribution shifts over time follow evolutionary patterns and has leveraged temporal information to capture such regularities. This perspective led to the Temporal Domain Generalization (TDG) [28; 40; 44; 2; 56; 59; 6], which models domain evolution along a temporal axis and seeks to generalize to future domains.
@@
A foundational modeling assumption in TDG is that domain evolution can be projected onto a single latent axis, typically instantiated as time. This simplification reduces domain generalization to a sequence extrapolation problem, enabling the use of well-established temporal modeling techniques. For instance, state-transition matrices simulate domain progression [44; 59]; recurrent neural networks capture temporal dependencies [2]; and neural ordinary differential equations model the continuous-time domain dynamics [6; 58]. While temporal formulation offers tractability, reducing domain evolution to a unidimensional trajectory imposes a fundamental limitation. In practice, distribution shifts rarely unfold along a single direction, but are instead continuously driven by multiple interacting fields such as climate conditions, socioeconomic status, infrastructure development, and population structure, each governing one aspect of the generative process. These factors jointly define a continuous variation space, over which domain shifts co-evolve smoothly. Collapsing such evolution onto a single temporal axis inevitably induces information loss and structural distortion.
@@
Figure 1: Illustration of continuous domain generalization. Real-world distributions are shaped by continuously varying factors. The observed domains provide sparse supervision over the joint model–variation space. Continuous domain generalization learns a predictive function over the space, enabling the model to generalize to any unseen domains along the underlying continuous structure.
@@
Formally, this paper introduces a problem termed Continuous Domain Generalization (CDG), which aims to generalize predictive models across domains that evolve over an arbitrary continuous space. This formulation captures a more general and realistic setting where domain variation arises continuously from multiple underlying factors. For instance, as shown in Fig. 1, in remote sensing, image semantics vary with seasonal climate and regional development level, which jointly shape land appearance: climate governs vegetation coverage and coloration, while development determines infrastructure density and land use. A collection of observed domains spanning combinations like highly urbanized areas in winter or rural zones in summer, which calibrates a prediction model space of domain variation. A new region with moderate development and autumn conditions naturally lies in this space. Its predictive model should emerge through evolution along the continuous variation field, yielding predictions that are both semantically coherent and structurally grounded. Such continuous domain generalization arises naturally in a wide range of applications. In healthcare, patient distributions may vary continuously with age, blood pressure/glucose, cholesterol levels, and body mass index. In urban analytics, regional traffic patterns evolve jointly with temporal variables and economic activity density. In visual recognition, image domains shift with illumination, viewpoint, and camera parameters like focal length. All of which continuously alter the underlying decision function.
@@
Continuous domain generalization introduces a novel and broadly applicable formulation. However, it poses three fundamental challenges: 1) Identifying domain variation to model evolution in general spaces. CDG requires characterizing how domain shifts over a high-dimensional space induce transitions in model behavior. However, this relationship remains theoretically unclear, with no principled constraints to ensure the coherent evolution of models across domains. 2) Structural indeterminacy necessitates a paradigm shift in modeling. Temporal domain evolution benefits from an inherent ordering, which enables the use of mature sequential modeling techniques. In contrast, general continuous spaces lack such priors, impeding the recovery of global structure and making it significantly more difficult to infer variation fields from sparse and irregular observations. 3) Leveraging under imperfect representation of domain variation. While CDG leverages readily available contextual variables as surrogate signals to guide generalization, these proxies are often noisy, incomplete, or misaligned with the true underlying variation. Such discrepancies introduce uncertainty into the learned parameter transitions and undermine generalization performance.
@@
To address all the challenges, we propose a generic continuous domain generalization framework, which extrapolates the model parameters in unseen domain using the observed domains in continuous space by transport operator techniques. We theoretically establish that optimal model parameters form a low-dimensional geometric manifold, providing a principled foundation for CDG task modeling. Building on this, we synergize Lie Group with a neural transport operator that enforces both geometric continuity and algebraic consistency. To ensure robust generalization under real-world imperfections, we incorporate a gating mechanism and a local chart-based strategy. Our framework jointly resolves the theoretical, structural, and practical difficulties inherent in CDG. Extensive experiments on both synthetic and real-world datasets, including remote sensing, scientific documents, and traffic forecasting, demonstrate that our method significantly outperforms existing baselines in generalization accuracy and robustness under descriptor imperfections.
@@ 2 Related works
Modeling Relationships Across Domains. Beyond treating domains as independent categories, a line of research seeks to model relationships or structures among domains. Before deep learning, early works modeled intermediate domain shifts via feature flows, such as geodesic flow kernels [18] and manifold alignment [20]. These methods relied on handcrafted features and linear subspace, limiting their applicability to modern deep, high-dimensional representations. Temporal Domain Generalization (TDG) [28; 40; 44; 2; 59; 6] extends the idea by assuming domains follow a sequential order, enabling recurrent or extrapolative generalization. However, these methods are restricted to the temporal axis and fail to generalize to continuous domains characterized by multi-dimensional, non-sequential variation, limiting their applicability in broader settings. Recent methods such as CIDA [51] and AdaGraph [37] incorporate domain index or metadata to model domain transitions. Despite their innovation, both operate in domain adaptation settings and require access to target domains during training. Overall, while these methods capture partial domain relationships, they are limited in scope and fail to capture the full structure of continuous domain variation.
Geometric and Algebraic Structures in Learning. Geometric deep learning aims to uncover the low-dimensional structures and regularities inherent in real-world tasks, formulating them through unified geometric and algebraic principles [5]. Mathematical tools such as manifolds, groups, topology, and vector fields provide a rigorous basis for encoding a wide range of physical regularities like invariance, equivariance, symmetry, locality, continuity, and conservation. These tools have led to significant advances in scientific machine learning and physics-informed modeling [45; 21; 8; 16; 34; 24], which demands reliable generalization under partial observability. Building on this foundation, we introduce a structural perspective on modeling for continuous domain generalization.
@@ 3 Problem definition
CDG addresses the task of generalizing predictive models across domains whose data distributions evolve continuously with respect to underlying latent factors. Formally, let $\{\mathcal{D}_{i}\}_{i=1}^{N}$ be a collection of domains, where each domain $\mathcal{D}_{i}$ consists of a set of input-output pairs drawn from a distribution over $\mathcal{X}\times\mathcal{Y}$, where $\mathcal{X}$ and $\mathcal{Y}$ denote the input and output spaces, respectively. Besides, each domain is associated with a descriptor $z_{i}\in\mathbb{R}^{d}$, which lies in a continuous descriptor space $\mathcal{Z}$ capturing domain-specific attributes such as spatial, temporal, demographic, economic, or environmental factors. We assume that the underlying data distribution in each domain is implicitly governed by its descriptor, meaning that the conditional distribution $P(Y_{i}\mid X_{i},z_{i})$, with $X_{i}$ and $Y_{i}$ as the datasets for inputs and targets under domain $\mathcal{D}_{i}$, respectively, evolves smoothly over $\mathcal{Z}$ as $z_{i}$ varies. This implies that small perturbations in $z_{i}$, either along individual coordinates or jointly, induce small changes in the conditional distribution.
@@
The goal of CDG is, during training, we are provided with a set of $N$ domains $\{\mathcal{D}_{1},\mathcal{D}_{2},\ldots,\mathcal{D}_{N}\}$, associated with their descriptors $\{z_{1},z_{2},\ldots,z_{N}\}$. For each domain $\mathcal{D}_{i}$, we learn a predictive model $g(\cdot;\theta_{i}):\mathcal{X}\rightarrow\mathcal{Y}$, where $\theta_{i}\in\mathbb{R}^{D}$ denotes the model parameters learned from $\mathcal{D}_{i}$. We want to learn the co-variations between domain descriptors $\{z_{1},z_{2},\ldots,z_{N}\}$ and model parameters $\{\theta_{1},\theta_{2},\ldots,\theta_{N}\}$, and to infer the parameters $\theta_{s}$ for an unseen domain $\mathcal{D}_{s}$ by giving a new descriptor $z_{s}\notin\{z_{1},\ldots,z_{N}\}$. This yields a predictive model $g(\cdot;\theta_{s})$ that generalizes to arbitrary domains across the continuous descriptor space $\mathcal{Z}$.
This continuity assumption is well-motivated by real-world phenomena, where distribution shifts are predominantly driven by continuously evolving processes. These dynamics typically induce smooth and incremental changes in the underlying data distributions across domains. Such regularities are widely observed in diverse applications and provide a tractable foundation for CDG. However, generalizing predictive models over such continuous spaces poses significant challenges. The existing temporal domain generalization method imposes a global ordering. In contrast, CDG operates over a general space, where no such canonical direction or ordering exists. This introduces structural indeterminacy and complicates both learning and generalization. In addition, standard domain adaptation methods rely on access to target domain input data during training, which CDG explicitly disallows and only leverages domain-level publicly available or easily accessible descriptors.
@@ 4 Methodology
In this section, we present our framework for continuous domain generalization via structured model transport. Specifically, we first identify the geometric structure of model parameters across domains and formally prove that the parameter set forms an embedded submanifold (Section 4.1). We then propose a Neural Lie Transport Operator (NeuralLTO) to enable structured, equivariant parameter transitions across domains, grounded in Lie group theory and respecting both geometric and algebraic constraints (Section 4.2). Finally, we introduce practical mechanisms to handle imperfections in the descriptor space, including degeneracy and incompleteness, through gating and local chart-based modeling (Section 4.3).
@@ 4.1 Identifying the Parameter Manifold of Continuous Domains
In this subsection, we reveal a coherent geometric structure underlying the optimal model parameters across continuous domains. Preserving this structure is fundamental for continuous generalization.
@@
In continuous domain generalization, the function mapping domain descriptors to their corresponding predictive model parameters, $\theta(z):\mathcal{Z}\to\Theta$, is continuous. Moreover, if the descriptor $z\in\mathbb{R}^{d}$ provides a complete and non-degenerate representation, then the image set  [^1] The image set of a function refers to the set of all values it outputs. :=$\mathcal{M}:=\{\theta(z)\mid z\in\mathcal{Z}\}$ is a $d$-dimensional embedded submanifold of $\mathbb{R}^{D}$.
@@
If $z$ provides a complete and non-degenerate representation, then the Jacobian $J_{\theta}(z)\in\mathbb{R}^{D\times d}$ is of full rank $d$, and the mapping $z\mapsto P(Y|X;z)$ is injective. Since the model architecture $g$ is fixed, $\theta(z)$ is uniquely determined by the conditional distribution, and hence the mapping $z\mapsto\theta(z)$ is also injective. Then by the Constant Rank Theorem, the image
:=$\mathcal{M}:=\{\theta(z)\mid z\in\mathcal{Z}\}$ (3)
is a $d$-dimensional embedded submanifold of $\mathbb{R}^{D}$. ∎
@@
The manifold structure regularizes the parameter space and constrains the solution to a semantically valid subset, reducing the searching complexity in $\mathbb{R}^{D}$.
@@
The manifold structure enables principled algebraic and geometric operations on the function of $g$, facilitating interpretable and controllable generalization.
4.2 Neural Domain Transport Operator under Structured Constraints
Eq. (2) reveals that a differential equation governs the evolution of model parameters. In the temporal case of domain generalization, prior work [6] models this evolution using NeuralODEs [9]. A natural extension is to lift the ODE to a PDE framework for multi-dimensional systems. However, this introduces substantial challenges: (1) Classical numerical PDE solvers, such as finite difference or finite element methods, require known values of $\theta(z)$ over a structured grid, but these $\theta(z)$ need to be jointly learned. (2) Modern neural PDE solvers like PINNs [45] rely on the symbolic form of the governing equations to construct residual losses. However, the governing PDE that describes the evolution of variation is unknown in CDG. (3) The observed domains are sparsely and irregularly distributed, precluding natural ordering or integration paths, making it infeasible to reduce the PDE to a set of tractable ODEs.
Structured Constrained Domain Transport
Rather than solving the PDE or explicitly regressing the parameter field $\theta(z)$, we propose to learn a structured transport operator $\mathcal{T}:\Theta\times\mathcal{Z}\times\mathcal{Z}\rightarrow\Theta$, which maps parameters at one descriptor $z_{i}$ to another target descriptor $z_{j}$. This structured operator circumvents limitations of classical and neural PDE solvers by directly learning pairwise parameter transitions. Specifically, given $\theta(z_{i})$ at $z_{i}$ and the target point $z_{j}$, the operator produces

### v2

@@ Abstract
Real-world data distributions often shift continuously across multiple latent factors such as time, geography, and socioeconomic contexts. However, existing domain generalization approaches typically treat domains as discrete or as evolving along a single axis (e.g., time). This oversimplification fails to capture the complex, multidimensional nature of real-world variation. This paper introduces the task of Continuous Domain Generalization (CDG), which aims to generalize predictive models to unseen domains defined by arbitrary combinations of continuous variations. We present a principled framework grounded in geometric and algebraic theories, showing that optimal model parameters across domains lie on a low-dimensional manifold. To model this structure, we propose a Neural Lie Transport Operator (NeuralLio), which enables structure-preserving parameter transitions by enforcing geometric continuity and algebraic consistency. To handle noisy or incomplete domain variation descriptors, we introduce a gating mechanism to suppress irrelevant dimensions and a local chart-based strategy for robust generalization. Extensive experiments on synthetic and real-world datasets, including remote sensing, scientific documents, and traffic forecasting, demonstrate that our method significantly outperforms existing baselines in both generalization accuracy and robustness. Code is available at: https://github.com/Zekun-Cai/NeuralLio.
footnotetext: * Corresponding author
@@ 1 Introduction
Distribution shift refers to changes in data distributions between training and deployment environments, where the input-output relationship learned during training no longer holds at test time. This discrepancy fundamentally compromises the reliability of learned models and motivates the study of Domain Generalization (DG) [52; 49; 23; 14; 43], which aims to train models on multiple source domains that generalize to unseen domains without accessing target-domain data. While early DG methods treat domains as independent entities, recent work recognizes that distribution shifts over time follow evolutionary patterns and thus leverages temporal information to capture such regularities. This perspective has led to the Temporal Domain Generalization (TDG) [28; 40; 44; 2; 56; 59; 6], which models domain evolution along a temporal axis and seeks to generalize to future domains.
@@
A foundational modeling assumption in TDG is that domain evolution can be projected onto a single latent axis, typically instantiated as time. This simplification reduces domain generalization to a model extrapolation problem, enabling the use of well-established sequence modeling techniques. For instance, state transition matrices simulate domain progression [44; 59]; recurrent neural networks capture temporal dependencies [2]; and ordinary differential equations model the continuous-time domain dynamics [6; 58]. While the temporal formulation offers tractability, reducing domain evolution to a unidimensional sequence imposes a fundamental limitation. In practice, distribution shifts rarely unfold along a single direction but are instead continuously driven by multiple interacting factors such as climate conditions, socioeconomic status, infrastructure development, and population structure, each governing one aspect of the generative process. These factors jointly define a continuous variation space, over which distribution shifts co-evolve smoothly. Collapsing such evolution onto a single temporal axis inevitably induces information loss and structural distortion.
@@
Figure 1: Illustration of continuous domain generalization. Real-world distributions are shaped by continuously varying factors. The observed domains provide sparse supervision over the joint variation-model space. Continuous domain generalization studies how predictive functions evolve over this space, enabling the model to generalize to unseen domains across the underlying continuous fields.
@@
Formally, this paper introduces a problem termed Continuous Domain Generalization (CDG), which aims to generalize predictive models across domains that evolve over an arbitrary continuous space. This formulation captures a more general and realistic setting where domain variation arises continuously from multiple underlying factors. For instance, as shown in Fig. 1, in remote sensing, image semantics vary with seasonal climate and regional development level, which jointly shape land appearance: climate governs vegetation coverage and coloration, and development determines infrastructure density and land use. A collection of observed domains, spanning combinations such as urban areas in winter or rural areas in summer, provides supervision to calibrate predictive models across the variation space. A new region with moderate development and autumn conditions naturally resides within this space, and its corresponding predictive model should emerge through evolution along the continuous variation field, yielding predictions that are both semantically coherent and structurally grounded. This mechanism similarly governs many real-world applications. In healthcare, patient distributions may vary continuously with age, blood pressure, glucose, cholesterol, and body mass index. In urban analytics, regional traffic patterns evolve jointly with time and economic activity density. In visual recognition, image domains shift with illumination, viewpoint, and camera parameters such as focal length. All of these factors continuously alter the underlying decision function.
@@
Continuous domain generalization introduces a novel and broadly applicable formulation. However, it poses three fundamental challenges: 1) Identifying domain variation to model evolution in general spaces. Continuous domain generalization requires characterizing how domain variation over a high-dimensional space induces associated transitions in model behavior. However, this relationship remains theoretically unclear, with no principled constraints to ensure the coherent evolution of models across domains. 2) Modeling without inherent structural priors of domains. Temporal domain generalization benefits from the chronological structure of domains, which enables the use of mature sequential modeling techniques. In contrast, general continuous space lacks such inherent topology, making it considerably challenging to recover the global structure of domains and to infer continuous variation fields from sparse and irregular observations. 3) Learning under imperfect representations of domain variation. While continuous domain generalization leverages readily available contextual variables as descriptor signals to guide generalization, these proxies are often noisy, incomplete, or misaligned with the true underlying variation. Such discrepancies introduce uncertainty into the learned parameter transitions and undermine generalization performance.
@@
To address these challenges, we propose a generic continuous domain generalization framework, which generalizes the model parameters to unseen domains using observed domains across continuous space by transport operator techniques. We theoretically show that the family of domain-optimal models constitutes a low-dimensional geometric manifold, providing a principled foundation for modeling continuous domain generalization. Building on this insight, we integrate Lie Group theory with a neural transport operator to enforce both geometric continuity and algebraic consistency. To ensure robust generalization under real-world imperfections, we incorporate a gating mechanism and a local chart-based strategy. Our framework jointly resolves the theoretical, structural, and practical challenges in continuous domain generalization. Extensive experiments on synthetic and real-world datasets, including remote sensing, scientific documents, and traffic forecasting, demonstrate that our method significantly outperforms existing baselines in both generalization accuracy and robustness.
@@ 2 Related Works
Modeling Relationships Across Domains. Beyond treating domains as independent categories, a line of research seeks to model relationships or structures among domains. Before deep learning, early works modeled intermediate domain shifts via feature flows, such as geodesic flow kernels [18] and manifold alignment [20]. These methods relied on handcrafted features and linear subspace, limiting their applicability to modern deep, high-dimensional representations. Temporal Domain Generalization (TDG) [28; 40; 44; 2; 59; 6] extends the idea by assuming domains follow a sequential order, enabling recurrent or extrapolative generalization. However, these methods are restricted to the temporal axis and fail to generalize to continuous domains characterized by multi-dimensional, non-sequential variation, limiting their applicability in broader settings. Other approaches leverage partial domain relationships based on domain indices. For instance, AdaGraph [37] constructs graphs over domain indices, while CIDA [51] defines domain relations through pairwise distances of domain indices. However, graphs are discrete, and distance-based formulations only induce a metric space, all of which lack the ability to express a global mapping over the domain variation space. Overall, while these methods capture partial domain relationships, they are limited in scope and fail to capture the full structure of continuous domain variation.
Geometric and Algebraic Structures in Learning. Geometric deep learning aims to uncover the low-dimensional structures and regularities underlying real-world data, formulating learning problems through unified geometric and algebraic principles [5]. It views data and models as entities living on structured spaces rather than in Euclidean coordinates, allowing learning to exploit intrinsic relationships. A broad spectrum of mathematical tools, including topology, geometry, group theory, graphs, and differential operators, provide rigorous formalisms to express physical and statistical regularities such as symmetry, invariance, equivariance, continuity, and conservation. These foundations have catalyzed major advances in scientific machine learning and physics-informed modeling [45; 21; 8; 16; 34; 24], enabling neural architectures to obey governing equations and preserve physical symmetries. Building on this foundation, we introduce a structural modeling perspective across domain-specific models for continuous domain generalization.
@@ 3 Problem Definition
CDG addresses the task of generalizing predictive models across domains whose data distributions evolve continuously with respect to underlying latent factors. Formally, let $\{\mathcal{D}_{i}\}_{i=1}^{N}$ be a collection of domains, where each domain $\mathcal{D}_{i}$ consists of a dataset $(X_{i},Y_{i})$ containing input–output pairs sampled from a distribution over $\mathcal{X}\times\mathcal{Y}$, with $\mathcal{X}$ and $\mathcal{Y}$ denoting the input and output spaces, respectively. In addition, each domain is associated with a descriptor $z_{i}\in\mathbb{R}^{d}$, with $z_{i}$ lying in a continuous descriptor space $\mathcal{Z}$ that captures domain-specific attributes such as spatial, temporal, demographic, economic, or environmental factors. We assume that the underlying data distribution of each domain is continuously governed by its descriptor, that is, the conditional distribution $P(Y\mid X,z_{i})$ evolves smoothly over $\mathcal{Z}$ as $z_{i}$ varies. Accordingly, small perturbations in $z_{i}$, either along individual coordinates or jointly, induce proportionally small changes in the conditional distribution.
@@
The goal of CDG is, during training, we are provided with a set of $N$ domains $\{\mathcal{D}_{1},\mathcal{D}_{2},\ldots,\mathcal{D}_{N}\}$, associated with their descriptors $\{z_{1},z_{2},\ldots,z_{N}\}$. For each domain $\mathcal{D}_{i}$, we learn a predictive model $g(\cdot;\theta_{i}):\mathcal{X}\rightarrow\mathcal{Y}$, where $\theta_{i}\in\mathbb{R}^{D}$ denotes the model parameters learned from $\mathcal{D}_{i}$. We want to learn the co-variations between domain descriptors $\{z_{1},z_{2},\ldots,z_{N}\}$ and model parameters $\{\theta_{1},\theta_{2},\ldots,\theta_{N}\}$, and to infer the parameters $\theta_{s}$ corresponding to an unseen domain $\mathcal{D}_{s}$ given a new descriptor $z_{s}\notin\{z_{1},\ldots,z_{N}\}$. This enables the learned framework to generalize to arbitrary domains across the descriptor space $\mathcal{Z}$, yielding a predictive model $g(\cdot;\theta_{s})$ for any given descriptor $z_{s}$.
CDG assumes that data distributions evolve continuously, consistent with the continuum hypothesis in physics and engineering, which posits that macroscopic systems change smoothly with respect to their underlying causal factors. Such smooth evolution is widely observed in physical, biological, and socioeconomic systems. Consequently, modeling domain evolution as a continuous process offers a principled and tractable abstraction of real-world dynamics. Nevertheless, generalizing predictive models over such continuous spaces poses significant challenges. First, domain variation and model behavior are often entangled, making it difficult to establish stable correspondences. Second, the absence of inherent chronological organization leaves no on-hand structural guidance for modeling cross-domain relationships. Third, domain-level information is typically noisy, incomplete, or misaligned, introducing uncertainty that undermines reliable generalization. In the following sections, we tackle these challenges sequentially.
@@ 4 Methodology
In this section, we present our framework for continuous domain generalization through structural model transport. Specifically, we first identify the geometric structure underlying model parameters across domains, and formally prove that the collection of domain-wise parameters forms an embedded submanifold. (Section 4.1). This establishes the geometric foundation for representing model evolution as a continuous mapping on the manifold. We then derive the necessary structural constraints that should be satisfied for model evolution, and propose a Neural Lie Transport Operator (NeuralLio) grounded in Lie Group theory to enable equivariant parameter transitions across domains (Section 4.2). Finally, we handle imperfections in the descriptor space, including degeneracy and incompleteness, using gating and local chart-based modeling (Section 4.3).
@@ 4.1 Identifying the Parameter Manifold of Continuous Domains
This subsection provides a theoretical perspective on the geometric structure of optimal model parameters across continuous domains, revealing a smooth and coherent mapping from domain descriptors to model parameters. Preserving this structure is fundamental for continuous generalization.
@@
In continuous domain generalization, the function mapping domain descriptors to their corresponding predictive model parameters, $\theta(z):\mathcal{Z}\to\Theta$, is continuous. Moreover, let $\mathcal{Z}\subseteq\mathbb{R}^{d}$ denote a theoretical descriptor space that provides a complete and non-degenerate representation of all latent factors governing domain variation, then the image set  [^1] The image set of a function refers to the set of all values it outputs. $\mathcal{M}:=\{\theta(z)\mid z\in\mathcal{Z}\}$ forms a $d$-dimensional embedded submanifold of $\mathbb{R}^{D}$.
@@
If the descriptor space $\mathcal{Z}$ provides a complete and non-degenerate representation of all latent factors governing domain variation  [^2] Such a descriptor space theoretically exists; in practice, observed descriptor serves as approximations without affecting the validity of the theorem., then the Jacobian $J_{z}\theta(z)\in\mathbb{R}^{D\times d}$ is of full rank $d$, and the mapping $z\mapsto P(Y|X;z)$ is injective. Since the model architecture $g$ is fixed, and each domain-specific parameter $\theta(z)$ is uniquely determined by the corresponding conditional distribution $P(Y|X;z)$, the mapping $z\mapsto\theta(z)$ is also injective. Then, by the Constant Rank Theorem, the image set
$\mathcal{M}:=\{\,\theta(z)\mid z\in\mathcal{Z}\,\}$ (3)
forms a $d$-dimensional embedded submanifold of $\mathbb{R}^{D}$.
@@
The manifold structure regularizes the parameter space by constraining optimization to a geometrically valid subset, reducing the searching complexity in $\mathbb{R}^{D}$.
@@
The manifold structure enables principled geometric and algebraic operations over model functions, facilitating interpretable and controllable generalization.
4.2 Neural Domain Transport Operator under Structural Constraints
Eq. (2) reveals that a differential equation governs the evolution of model parameters. In the temporal case where the domain descriptor is one-dimensional, prior work [6] models such evolution with NeuralODEs [9]. A natural extension is to lift the ODE to a PDE framework for multi-dimensional cases. However, this introduces substantial challenges: (1) Classical numerical PDE solvers, such as finite difference or finite element methods, require known values of $\theta(z)$ on a structured grid to specify boundary or initial conditions. However, in continuous domain generalization, the $\theta(z)$ for each training domain is not directly known and must be learned through iterative optimization, leaving such conditions undefined. (2) The training domains are sparsely and irregularly distributed, precluding natural ordering or integration paths, making it infeasible to reduce the PDE to a set of ODEs along any direction. (3) Modern neural PDE solvers, such as Physics-Informed Neural Networks [45], rely on symbolic PDE formulations to define residual losses, yet the governing equations describing distributional evolution remain unknown in continuous domain generalization.
Structural Constrained Domain Transport
Rather than solving the PDE or explicitly regressing the parameter field $\theta(z)$, we propose to learn a structure-preserving transport operator $\mathcal{T}:\Theta\times\mathcal{Z}\times\mathcal{Z}\rightarrow\Theta$, which maps parameters at one descriptor $z_{i}$ to another target descriptor $z_{j}$. This structural operator circumvents limitations of classical and neural PDE solvers by directly learning pairwise parameter transitions. Specifically, given $\theta(z_{i})$ at $z_{i}$ and the target point $z_{j}$, the operator produces